For this inverted T shape, what is the position of its centroid relative to the bottom edge of the shape? * 2 * 2 * 2 →
Answer #1

2B The Area of the rectangle ABCH = (542)=10 sq.unit =A(say) The Area of the rectangle DEFG=(2+2+2X2)=12 sq.unit =A2 (say) No